OpenStudy (anonymous):

What is the slope of the line? y = -2x - 8 Please help! :)

pooja195 (pooja195):

y=mx+b m=slope

OpenStudy (anonymous):

Where do I get the B?

pooja195 (pooja195):

b is the y intercept

pooja195 (pooja195):

We only need to worry about m the slope would be the number in front of the x .

pooja195 (pooja195):

Hence. y=mx+b m=slope

OpenStudy (anonymous):

But how do I solve the equation if I don't know what the slope is?

pooja195 (pooja195):

Your question asks what is the slope of the line y=mx+b <---slope intercept form m=slope b=y intercept there is no solving the slope is already given

pooja195 (pooja195):

Knowing that info m=?

OpenStudy (anonymous):

Thanks for your help. I am solving it now.

pooja195 (pooja195):

There is no solving needed though save yourself the extra time lol

OpenStudy (anonymous):

.. I'm sorry but you really aren't making it easy to understand what you're telling me.

pooja195 (pooja195):

We are given an equation y = -2x - 8 The equation is in the form of a slope intercept equation y=mx+b is the slope intercept form m is the slope b is the y intercept the number in front of x is the m which is the slope

pooja195 (pooja195):

What number (in the equation) is in front of the x?

OpenStudy (anonymous):

-2..

pooja195 (pooja195):

Good so our slope is -2 Do you understand?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but thanks anyway for trying

pooja195 (pooja195):

What do you need help with? What don't you understand? How may I help you better? :-)

OpenStudy (anonymous):

Nothing really, I'll figure the rest out on my own. Thank you

pooja195 (pooja195):

yw